### v2

@@ 2.2 Universal approximation theorem
Hua et al. (2023) [16] consider a general, possibly nonlinear, operator $\mathcal{J}$ between two distinct Hilbert spaces.
@@
In this work, we restrict ourselves to the transfer and Koopman operators $\mathcal{L}$ associated with a dynamical system. Each of these operators map $L^{2}(X)$ into $L^{2}(X)$, so only a single subspace $V\subset L^{2}(X)$ is required to approximate both the observable $g$ and its image $\mathcal{L}(g)$. Consequently, in contrast to [16], one should instead use a single encoder. We state an approximation theorem in this alternate setting.
@@
Let $X\subset\mathbb{R}^{d}$ be compact and let $T:X\to X$ be the evolution dynamics on $X$. Let $\mathcal{L}:L^{2}(X)\rightarrow L^{2}(X)$ be continuous and denote either the Koopman or the transfer operator associated with $T$. Given $N\in\mathbb{N}$ and a family $\phi=(\phi_{1},\dots,\phi_{N})\subset L^{2}(X)$ define
@@
Then for each compact set $\mathcal{K}\subset L^{2}(X)$ and any $0<\epsilon<1$ there exist $N>0$, a neural basis $\phi=(\phi_{1},\phi_{2},\dots,\phi_{N}):X\rightarrow\mathbb{R}^{N}$ and a linear map $\mathcal{G}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ such that the operator ^$\hat{\mathcal{L}}\coloneq\mathcal{R}_{\phi}\circ\mathcal{G}\circ\mathcal{P}_{\phi}$ satisfies
@@
Examples of compact sets $\mathcal{K}\in L^{2}(X)$ with $X\subset\mathbb{R}^{d}$, include (i) the closed unit ball of any finite-dimensional subspace in $L^{2}(X)$ and (ii) the closed unit ball of $H^{1}(X)$, for $d\geq 2$ and $X$ a bounded $C^{1}$ domain (its boundary is locally a $C^{1}$ graph). The latter is compact in $L^{2}(X)$ by the Rellich–Kondrachov Embedding Theorem (e.g. Theorem 9.16 [6]).
Two convenient sufficient conditions that guarantee the continuity of $\mathcal{L}$ are:
@@
    $\mu$ almost everywhere. The Anosov map in Section 3.4 is a representative example with $\mu=$ Lebesgue.
@@
The proof proceeds almost exactly as in [16, Theorem 2.2], so we omit the repeated details and focus on the necessary modifications. We need only one neural basis $\{\phi_{1},\ldots,\phi_{N}\}\subset L^{2}(X)$ for both the input and output spaces. The universal approximation property [16, Theorem 2.1] ensures we can approximate each $g\in\mathcal{K}$ by an element $g^{N}\in\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$. Let $\mathcal{K}_{\mathcal{L}}:=\mathcal{K}\cup\mathcal{L}(\mathcal{K})$ and choose $N$ and $\phi$ by [16, Theorem 2.1] so that, writing $Q:=\mathcal{R}_{\phi}\mathcal{P}_{\phi}$,
@@
The main difference appears in the error decomposition. In [16], one bounds three terms to handle potentially different neural bases for input and output spaces, but here we merge everything into two terms. If ^$\hat{\mathcal{L}}\coloneq\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{G}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}$, then for $g\in\mathcal{K}$
@@
+$\quad+\Bigl{\|}\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{L}(g)-\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{G}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}(g)\Bigr{\|}_{L^{2}(X)}.$
@@
Everything else remains unchanged from [16, Theorem 2.2], and thus we conclude
@@ 3 Numerical examples
In this section, we illustrate the results obtained with our architecture through two representative examples. We consider the Perron–Frobenius operator in both examples, but the procedure is identical for training data generated by the Koopman operator. The first example considers a circle rotation, highlighting the recovery of spectral properties of the associated linear operator. The second example examines a nonlinearly perturbed version of Arnold’s cat map on the two-torus. This nonlinear map is highly anisotropic in state space and the eigenfunctions of the transfer operator are known to be distributions rather than functions [15, 9]. Through this very challenging example, we demonstrate how the learned basis functions can nevertheless successfully adapt to the system dynamics to construct an almost perfectly invariant subspace $V$.
@@ 3.1 Data generation
Both examples in this paper — the circle rotation and the perturbed cat map — are dynamical systems on periodic domains. We therefore require periodic observables, and we use random trigonometric polynomials to provide a rich class of training functions. To train the proposed architecture, we consider the intrinsic dimension $d^{\prime}=\dim\;X$ and generate a dataset of $D$ random trigonometric polynomial functions defined on a $d^{\prime}$-dimensional domain. Each such function is constructed as a finite linear combination of sinusoidal basis functions (sine and cosine terms) in each dimension, with randomly chosen coefficients. Formally, a general $d^{\prime}$-dimensional trigonometric polynomial of maximum order $K$ can be written as a linear combination of terms of the form $\prod_{j=1}^{d^{\prime}}\Psi_{j}(k_{j}x_{j})$, where each $\Psi_{j}(\cdot)$ is either $\sin(\cdot)$ or $\cos(\cdot)$, and each integer frequency $k_{j}$ ranges from 111 to $K$ for $\sin(\cdot)$ and from 0 to $K$ for $\cos(\cdot)$. In other words, no oscillatory component in any dimension exceeds frequency $K$. The coefficients in the linear combination are real numbers drawn at random uniformly in [1,1][-1,1][ - 1 , 1 ]. Note that the inclusion of $k_{j}=0$ allows constant (zero-frequency) components in each dimension as well. For example, in $d^{\prime}=2$ dimensions (with coordinates $x_{1}$ and $x_{2}$), such trigonometric polynomials take the form:
$f(x_{1},x_{2})=$ $c_{1}$ $\sin(k_{1}x_{1})\sin(k_{2}x_{2})+c_{2}\sin(k_{1}x_{1})\cos(k_{2}x_{2})$
$+$ $c_{3}\cos(k_{1}x_{1})\sin(k_{2}x_{2})+c_{4}\cos(k_{1}x_{1})\cos(k_{2}x_{2}),$
with $k_{1},k_{2}\in\{1,\ldots,K\}$ for the sine terms and $k_{1},k_{2}\in\{0,1,\ldots,K\}$ for the cosine terms. In the general $d^{\prime}$-dimensional case, each term is a product of trigonometric functions, one for each of the $d^{\prime}$ coordinates, with frequencies up to $K$ in each coordinate. Such functions contain a rich variety of mixed-frequency components across all dimensions.
Once the $D$ random functions ${f_{j}}$ are generated as described above, we apply $\mathcal{L}$ to each of them to obtain the corresponding output functions $\mathcal{L}f_{j}$. This yields a training set of $D$ input-output pairs $\{f_{j};\mathcal{L}f_{j}\}_{j=1}^{D}$. In our training, $[f_{j}(x_{1}),f_{j}(x_{2}),\ldots,f_{j}(x_{n})]^{\top}$ serves as an input function sampled on the discretization $\{x_{i}\}_{i=1,\dots,n}$, and $[\mathcal{L}f_{j}(x_{1}),\allowbreak\mathcal{L}f_{j}(x_{2}),\ldots,\mathcal{L}f_{j}(x_{n})]^{\top}$ is treated as the target output. By using trigonometric polynomials as input functions, we ensure that the action of $\mathcal{L}$ is well defined and smooth on these inputs. This property simplifies the generation of training data when $\mathcal{L}f_{j}$ can be computed analytically, and preserves the continuity that is important for machine-learned approximations.
@@ 3.2 Examples setup
We train on collections of functions as described in Section 3.1. The validation and test sets are drawn from the same underlying class of functions as our training data. The parameters of each dataset are summarized in Table 1.
@@
Table 2: Architecture specifics for the circle rotation and perturbed cat map. We list the encoder ($\mathcal{E}$) and approximation networks ($\mathcal{G}$), their hidden layer architectures (e.g. 5×5125$\times$ 5125 × 512 denotes five hidden layers with 512 units each), and activation functions. The dashes indicate that there are no hidden layers for the network $\mathcal{G}$ since it is learning a linear map in $\mathbb{R}^{N}$.
Example Encoder ($\mathcal{E}$) Approximation ($\mathcal{G}$) N
Hidden layers Activ. Hidden layers Activ.
Circle rotation 5 $\times$ 512 ReLU - Identity 19
Perturbed cat map 5 $\times$ 2048 ReLU - Identity 225
@@ 3.3 Transfer operator for a circle rotation
Figure 2: Embedding of the circle‑rotation map $T(\theta)=\theta+\alpha\pmod{2\pi}$ from the interval $[0,2\pi)$ into the unit circle $S^{1}\subset\mathbb{R}^{2}$. Each angle $\theta$ is sent to its Cartesian coordinates $(\cos\theta,\sin\theta)$, removing the discontinuity that would occur at $\theta=0\equiv 2\pi$ in the interval picture. The initial point $\theta_{0}$ is chosen close to the right‑hand endpoint of the interval to highlight the jump that is resolved by the embedding.
After the learning routine, we obtain both a set of basis functions spanning a subspace invariant under the transfer operator and an estimator for its action. Fig. 3 displays how the learned estimator performs on unseen data for various sparsity scenarios.
@@
The approximately invariant subspace is spanned by the set of basis functions depicted in Fig. 4. By introducing a sparsity penalty, these functions gain increasingly localised support, as shown in Fig. 4.
Figure 4: Basis functions obtained under different sparsity penalties: each panel displays a distinct basis for a distinct sparsity penalty. In a single panel, the state space coordinates (the angles $\theta\in S^{1}$) are along the $x$-axis, and each row shows the values of a different basis function. Higher sparsity encourages localisation, which is reflected in more concentrated regions of support.
@@
Figure 5: Gram matrix of the filtered and normalised basis functions
@@
$g=\sum_{j=1}^{N}\beta_{j}\phi_{j},$
where $\beta=(\beta_{1},\dots,\beta_{N})^{T}\in\mathbb{R}^{N}$ is the coefficient vector.
@@
$\mathcal{P}_{\phi}(g)=\begin{pmatrix}\langle g,\phi_{1}\rangle\\ \vdots\\ \langle g,\phi_{N}\rangle\end{pmatrix}=\begin{pmatrix}\langle\sum_{j=1}^{N}\beta_{j}\phi_{j},\phi_{1}\rangle\\ \vdots\\ \langle\sum_{j=1}^{N}\beta_{j}\phi_{j},\phi_{N}\rangle\end{pmatrix}=M\beta,$
@@
The learned operator is composed as ^$\hat{\mathcal{L}}=\mathcal{R}_{\phi}\circ\mathcal{G}\circ\mathcal{P}_{\phi}$ where $\mathcal{G}$ is an $N\times N$ matrix representing the dynamics in coefficient space. When $\hat{\mathcal{L}}$ is applied to $g$ we obtain the following linear transformation:
^$\hat{\mathcal{L}}g=\mathcal{R}_{\phi}(\mathcal{G}(\mathcal{P}_{\phi}(g)))=\mathcal{R}_{\phi}(\mathcal{G}(M\beta))=\sum_{j=1}^{N}(\mathcal{G}M\beta)_{j}\phi_{j}.$
A function $g$ with coefficient vector $\beta$ is an eigenfunction of $\hat{\mathcal{L}}$ with eigenvalue $\lambda$ if and only if ^$\hat{\mathcal{L}}g=\lambda g$. This means
$\sum_{j=1}^{N}(\mathcal{G}M\beta)_{j}\phi_{j}=\lambda\sum_{j=1}^{N}\beta_{j}\phi_{j}.$
@@
(8) $\mathcal{G}M\beta=\lambda\beta.$
Equation (8) is a generalised eigenvalue problem of size $N\times N$. Solving it yields eigenvalue-eigenvector pairs $(\lambda_{k},\beta^{(k)})$ for $k=1,\dots,N$. The corresponding eigenfunctions are
$g_{k}=\sum_{j=1}^{N}\beta_{j}^{(k)}\phi_{j},\quad k=1,\dots,N.$
@@
which shows that we have eigenpairs $(\lambda_{k},\psi_{k}),k\in\mathbb{Z}$. Thus, the leading eigenvalues take the form
@@
Figure 6: Left: Eigenvalues of the Perron–Frobenius operator for a circle rotation by $\alpha=-1$ radians. Numerical eigenvalues (blue circles; leading pair corresponding to $k=\pm 1$ shown as red circles) coincide with the analytic spectrum (green crosses) Right: Magnitude versus argument of the same eigenvalues. The dashed green vertical lines mark angles that are positive integer multiples of $\alpha$, underscoring that the additional numerical eigenvalues are dynamically meaningful, but duplicate information from the lowest-order eigenvalues.
@@
Finally, we demonstrate the invariance of the learned subspace under the Perron–Frobenius operator. On the grid $\{x_{i}\}_{i=1,\dots,n}$ define
$\Phi=[\phi_{j}(x_{i})]_{i=1,\dots,n;\,j=1,\dots,N}\in\mathbb{R}^{n\times N},\qquad\Psi=[\mathcal{L}(\phi_{j})(x_{i})]_{i=1,\dots,n;\,j=1,\dots,N}\in\mathbb{R}^{n\times N},$
and set $r=\operatorname{rank}\Phi=\operatorname{rank}\Psi\leq N$. We wish to determine how close the $r$-dimensional column space of $\Phi$ is to the $r$-dimensional column space of $\Psi$, and we do this by computing principal angles between these two subspaces. Following [14, Algorithm 6.4.3] we form the QR decomposition
$\Phi=Q_{\Phi}R_{\Phi},\qquad\Psi=Q_{\Psi}R_{\Psi},$
with orthogonal $Q_{\Phi},Q_{\Psi}\in\mathbb{R}^{n\times n}$ and upper-triangular $R_{\Phi},R_{\Psi}\in\mathbb{R}^{n\times N}$. Denote by $Q_{\Phi}^{(r)}$ and $Q_{\Psi}^{(r)}$ the first $r$ columns of $Q_{\Phi}$ and $Q_{\Psi}$. If $0\leq\theta_{1}\leq\cdots\leq\theta_{r}$ are the principal angles between the subspaces spanned by $\Phi$ and $\Psi$, then
$\cos\theta_{k}=\sigma_{k}\!\bigl{[}(Q_{\Phi}^{(r)})^{\top}Q_{\Psi}^{(r)}\bigr{]},\qquad k=1,\ldots,r,$
To avoid numerical issues associated with the classical cosine-based computation for small angles and the sine-based approach for angles near $\frac{\pi}{2}$, we employed the combined technique [20] that uses the cosine-based method initially and switches to the sine-based approach for principal angles smaller than $\frac{\pi}{4}$. Note that the computation of principal angles is extremely sensitive to numerical precision. Therefore, we use higher precision floating-point arithmetic for these operations in both of our examples to ensure accurate results.
From these angles $\theta=(\theta_{1},\ldots,\theta_{N})$, we report three standard subspace‑distance metrics:
  1. 1.
    the geodesic distance $d_{\max}=\max_{k}\theta_{k}$,
  2. 2.
    the chordal distance $d_{\text{ch}}=\bigl{\|}\sin\theta\bigr{\|}_{2}$,
  3. 3.
    the projection distance $d_{\text{pr}}=\sin d_{\max}$.
Table 4: Subspace‐distance measures between $\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$ and $\operatorname{span}\{\mathcal{L}(\phi_{j})\}_{j=1}^{N}$
Metric $\beta_{2}=0$ $\beta_{2}=0.1$ $\beta_{2}=0.2$ $\beta_{2}=0.6$
Geodesic ($d_{\max}$) $3.983\times 10^{-2}$ $3.621\times 10^{-2}$ $3.869\times 10^{-2}$ $4.008\times 10^{-2}$
Chordal ($d_{\text{ch}}$) $3.982\times 10^{-2}$ $3.621\times 10^{-2}$ $3.868\times 10^{-2}$ $4.008\times 10^{-2}$
Projection ($d_{\text{pr}}$) $1.760\times 10^{-2}$ $1.906\times 10^{-2}$ $2.146\times 10^{-2}$ $1.732\times 10^{-2}$
Smaller values of the principal angles $\theta_{k}$ indicate that the two subspaces are closer, and hence the learned subspace approaches invariance under $\mathcal{L}$. Across all sparsity levels $\beta_{2}$, the reported distances remain on the order of $10^{-2}$, providing strong evidence of a learned subspace that is nearly invariant under the true operator $\mathcal{L}$.
@@
We close this example by discussing the choice of $N=19$ basis functions. Recall we trained $\hat{\mathcal{L}}$ using trigonometric polynomials up to order 9 (equivalently linear combinations of real Fourier modes up to order 9). As discussed above, each two-dimensional real subspace spanned by $\theta\mapsto\sin(k\theta)$ and $\theta\mapsto\cos(k\theta)$ is an eigenspace with eigenpair $\lambda_{\pm k}$. Thus, in this special situation of circle rotation dynamics, the input and output data lie in the same 19-dimensional (19=2×9+119=2$\times$ 9+119 = 2 × 9 + 1) subspace. Increasing the learned basis cardinality $N$ beyond 19 will not impact the prediction error, but the principal angles beyond the first 19 will increase because these additional dimensions will effectively be unable to be controlled with $9^{\rm th}$-order input/output data.
3.4 Transfer operator for a nonlinearly perturbed Arnold’s cat map

### v3

@@ 2.2 Universal approximation theorem
Hua et al. (2023) [HUA202321] consider a general, possibly nonlinear, operator $\mathcal{J}$ between two distinct Hilbert spaces.
@@
Each of the transfer and Koopman operators map $L^{2}(X)$ into $L^{2}(X)$, so only a single subspace $V\subset L^{2}(X)$ is required to approximate both the observable $g$ and its image $\mathcal{L}(g)$. Consequently, in contrast to [HUA202321], one should instead use a single encoder. We state an approximation theorem in this alternate setting.
@@
Let $X\subset\mathbb{R}^{d}$ be compact. Let $\mathcal{L}:L^{2}(X)\rightarrow L^{2}(X)$ be a continuous linear operator. Given $N\in\mathbb{N}$ and a family $\phi=\{\phi_{1},\dots,\phi_{N}\}\subset L^{2}(X)$ define
@@
Then for each compact set $\mathcal{K}\subset L^{2}(X)$ and any $0<\epsilon<1$ there exist $N>0$, a neural basis $\phi=\{\phi_{1},\phi_{2},\dots,\phi_{N}\}:X\rightarrow\mathbb{R}^{N}$ and a linear map $\mathcal{G}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ such that the operator $\hat{\mathcal{L}}\coloneq\mathcal{R}_{\phi}\circ\mathcal{G}\circ\mathcal{P}_{\phi}$ satisfies
@@
Although Theorem 2.1 applies to any continuous linear operator $\mathcal{L}:L^{2}(X)\rightarrow L^{2}(X)$, we will focus on the transfer and Koopman operators.
@@
Examples of compact sets $\mathcal{K}\subset L^{2}(X)$ with $X\subset\mathbb{R}^{d}$, include (i) the closed unit ball of any finite-dimensional subspace in $L^{2}(X)$ and (ii) the closed unit ball of $H^{1}(X)$, for $d\geq 2$ and $X$ a bounded $C^{1}$ domain (its boundary is locally a $C^{1}$ graph). The latter is compact in $L^{2}(X)$ by the Rellich–Kondrachov Embedding Theorem (e.g. Theorem 9.16 [brezis]).
In the context of the Koopman and transfer operators associated with a dynamical system $T:X\rightarrow X$, two convenient sufficient conditions that guarantee the continuity of $\mathcal{L}$ are:
@@
    $\mu$ almost everywhere. The Anosov map in Section 3.4.1 is a representative example with $\mu=$ Lebesgue.
@@
[…]
@@
[…]
@@
$\qquad+\Bigl\|\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{L}(g)-\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{G}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}(g)\Bigr\|_{L^{2}(X)}.$
@@
Everything else remains unchanged from [HUA202321, Theorem 2.2], and thus we conclude
@@ 3 Numerical examples
In this section, we illustrate the results obtained with our architecture through three representative examples of increasing complexity. We consider the Perron–Frobenius operator in all examples, but the procedure is identical for training data generated by the Koopman operator.
The first example considers a circle rotation, highlighting the recovery of meaningful eigenvalues and eigenfunctions of the transfer operator. The remaining two examples examine nonlinear Anosov diffeomorphisms on the two-torus: a weakly nonlinear perturbation of Arnold’s cat map, and a strongly nonlinear variant. These hyperbolic maps are highly anisotropic and the eigenfunctions of their transfer operators are known to be distributions rather than functions [GL06]. Through these challenging examples, we demonstrate how the learned basis functions adapt to the system dynamics, providing efficient and accurate recovery of spectral properties such as the SRB measure.
@@ 3.1 Data generation
All examples in this paper — the circle rotation and the two Anosov maps — are dynamical systems on periodic domains. We therefore require periodic observables, and we use random trigonometric polynomials to provide a rich class of training functions. To train the proposed architecture, we consider the intrinsic dimension $d^{\prime}=\dim\;X$ and generate a dataset of $D$ random trigonometric polynomial functions defined on a $d^{\prime}$-dimensional domain. Each such function is constructed as a finite linear combination of sinusoidal basis functions (sine and cosine terms) in each dimension, with randomly chosen coefficients. Formally, a general $d^{\prime}$-dimensional trigonometric polynomial of maximum order $K$ can be written as a linear combination of terms of the form $\prod_{j=1}^{d^{\prime}}\Psi_{j}(k_{j}x^{(j)})$, where each $\Psi_{j}(\cdot)$ is either $\sin(\cdot)$ or $\cos(\cdot)$, and each integer frequency $k_{j}$ ranges from $1$ to $K$ for $\sin(\cdot)$ and from 0 to $K$ for $\cos(\cdot)$. In other words, no oscillatory component in any dimension exceeds frequency $K$. The coefficients in the linear combination are real numbers drawn at random uniformly in [1,1][-1,1]. Note that the inclusion of $k_{j}=0$ allows constant (zero-frequency) components in each dimension as well. For example, in $d^{\prime}=2$ dimensions (with coordinates $x^{(1)}$ and $x^{(2)}$), such trigonometric polynomials are linear combinations of $\sin(k_{1}x^{(1)})\sin(k_{2}x^{(2)}),\sin(k_{1}x^{(1)})\cos(k_{2}x^{(2)}),\cos(k_{1}x^{(1)})\sin(k_{2}x^{(2)})$, and $\cos(k_{1}x^{(1)})\cos(k_{2}x^{(2)}),$ with $k_{1},k_{2}\in\{1,\ldots,K\}$ for the sine terms and $k_{1},k_{2}\in\{0,1,\ldots,K\}$ for the cosine terms. In the general $d^{\prime}$-dimensional case, each term is a product of trigonometric functions, one for each of the $d^{\prime}$ coordinates, with frequencies up to $K$ in each coordinate. Such functions contain a rich variety of mixed-frequency components across all dimensions.
Once the $D$ random functions $f_{j}$} are generated as described above, we apply $\mathcal{L}$ to each of them to obtain the corresponding output functions $\mathcal{L}f_{j}$. This yields a training set of $D$ input-output pairs $\{f_{j};\mathcal{L}f_{j}\}_{j=1}^{D}$. If the reconstruction loss $\mathcal{E}_{3}$ is used with index $k>1$, we additionally use the $k$-step iterates $\mathcal{L}^{k}f_{j}$. In our training, $f_{j}(x_{1}),f_{j}(x_{2}),\ldots,f_{j}(x_{n})]^{\top}$ serves as an input function evaluated on the discretization $\{x_{i}\}_{i=1,\dots,n}$, and $\mathcal{L}f_{j}(x_{1}),\allowbreak\mathcal{L}f_{j}(x_{2}),\ldots,\mathcal{L}f_{j}(x_{n})]^{\top}$ is treated as the target output. By using trigonometric polynomials as input functions, we ensure that the action of $\mathcal{L}$ is well defined and smooth on these inputs. This property simplifies the generation of training data when $\mathcal{L}f_{j}$ can be computed analytically, and preserves the continuity that is important for machine-learned approximations.
@@ 3.2 Examples setup
We train on collections of functions as described in Section 3.1. In our examples, the discretizations $\{x_{i}\}_{i=1,\dots,n}$ lie on a uniform grid in the domain, but we could also have chosen a discretization based on scattered points. The validation and test sets are drawn from the same underlying class of functions as our training data. The parameters of each dataset are summarised in Table 1.
@@
Table 2: Architecture specifics for the circle rotation and nonlinear cat map. We list the encoder ($\mathcal{E}$) hidden layer architecture (e.g. $5\times 512$ denotes five hidden layers with 512 units each), activation functions, total parameter count, and training time. The approximation network $\mathcal{G}$ is a linear map (i.e. a matrix) acting on $\mathbb{R}^{N}$ for all examples.
Example Encoder ($\mathcal{E}$) NN Params. Time
Layers Act.
Circle rotation 5 $\times$ 512 ReLU 19 1.06M 2.25 min
Nonlinear cat map 5 $\times$ 2048 ReLU 324 17.6M 17.65 min
Conj. nonlinear cat map 5 $\times$ 2048 ReLU 676 18.6M 37.13 min
All experiments were performed on an A100 NVIDIA GPU, and the model was implemented using PyTorch. We trained for 10,000 epochs on the circle rotation and 4,500 epochs on each Anosov map. We briefly comment on sensitivity to hyperparameters. As shown in Table 2, more complex dynamics may require a network with slightly larger capacity (e.g. wider hidden layers). However, the dominant factor affecting performance is the number of basis functions $N$, which must grow with the complexity of the dynamics to adequately capture the operator dynamics. Regarding the multi-step reconstruction loss $\mathcal{E}_{3}$, we observe that the choice of index $k$ should reflect the anisotropy of the dynamics. For systems with more irregular structures in certain directions — such as Anosov diffeomorphisms, whose eigendistributions are smooth along unstable manifolds but highly irregular along stable ones — larger values of $k$ are beneficial. This is because higher iterates $\mathcal{L}^{k}(g)$ increasingly expose the anisotropic geometry of the dynamics, encouraging the learned basis to adapt to these irregular directions. Across our three examples, we used $k=0$ for the circle rotation and $k=2$ for the two Anosov maps.
@@ 3.3 Transfer operator for a circle rotation
Figure 2: Embedding of the circle‑rotation map $T(\theta)=\theta+\alpha\pmod{2\pi}$ from the interval $0,2\pi)$ into the unit circle $S^{1}\subset\mathbb{R}^{2}$. Each angle $\theta$ is sent to its Cartesian coordinates $(\cos\theta,\sin\theta)$, removing the discontinuity that would occur at $\theta=0\equiv 2\pi$ in the interval picture. The initial point $\theta_{0}$ is chosen close to the right‑hand endpoint of the interval to highlight the jump that is resolved by the embedding.
After the learning routine, we obtain both a set of basis functions spanning a subspace well-suited for spectral approximation and an estimator $\hat{\mathcal{L}}$ for the transfer operator’s action. Fig. 3 displays how $\hat{\mathcal{L}}$ performs on unseen data for various sparsity scenarios.
@@
The learned subspace is spanned by the set of basis functions depicted in Fig. 4. By introducing a sparsity penalty, these functions gain increasingly localised support, as shown in Fig. 4.
Figure 4: $L^{2}$-normalised basis functions obtained under different sparsity penalties: each panel displays a distinct basis for a distinct sparsity penalty. In a single panel, the state space coordinates (the angles $\theta\in S^{1}$) are along the $x$-axis, and each row shows the values of a different basis function. Higher sparsity encourages localisation, which is reflected in more concentrated regions of support.
@@
Figure 5: Gram matrix of the normalised basis functions
@@
$g=\sum_{j=1}^{N}\xi_{j}\phi_{j},$
where $\xi=(\xi_{1},\dots,\xi_{N})^{T}\in\mathbb{R}^{N}$ is the coefficient vector.
@@
$\mathcal{P}_{\phi}(g)=\begin{pmatrix}\langle g,\phi_{1}\rangle\\ \vdots\\ \langle g,\phi_{N}\rangle\end{pmatrix}=\begin{pmatrix}\langle\sum_{j=1}^{N}\xi_{j}\phi_{j},\phi_{1}\rangle\\ \vdots\\ \langle\sum_{j=1}^{N}\xi_{j}\phi_{j},\phi_{N}\rangle\end{pmatrix}=M\xi,$
@@
The learned operator is $\hat{\mathcal{L}}=\mathcal{R}_{\phi}\circ\mathcal{G}\circ\mathcal{P}_{\phi}$ where $\mathcal{G}$ is an $N\times N$ matrix representing the dynamics in coefficient space. When $\hat{\mathcal{L}}$ is applied to $g$ we obtain the following linear transformation:
$\hat{\mathcal{L}}g=\mathcal{R}_{\phi}(\mathcal{G}(\mathcal{P}_{\phi}(g)))=\mathcal{R}_{\phi}(\mathcal{G}(M\xi))=\sum_{j=1}^{N}(\mathcal{G}M\xi)_{j}\phi_{j}.$
A function $g$ with coefficient vector $\xi$ is an eigenfunction of $\hat{\mathcal{L}}$ with eigenvalue $\lambda$ if and only if $\hat{\mathcal{L}}g=\lambda g$. This means
$\sum_{j=1}^{N}(\mathcal{G}M\xi)_{j}\phi_{j}=\lambda\sum_{j=1}^{N}\xi_{j}\phi_{j}.$
@@
(9) $\mathcal{G}M\xi=\lambda\xi.$
Equation (9) is an eigenproblem of size $N\times N$. Solving it yields eigenvalue-eigenvector pairs $(\lambda_{k},\xi^{(k)})$ for $k=1,\dots,N$. The corresponding eigenfunctions are
$g_{k}=\sum_{j=1}^{N}\xi_{j}^{(k)}\phi_{j},\quad k=1,\dots,N.$
@@
which shows that we have eigenpairs $(\lambda_{k},\psi_{k}),k\in\mathbb{Z}$. Furthermore, for an irrational rotation, these are the only eigenpairs: the Fourier modes form an orthonormal basis for $L^{2}(S^{1})$, and the $2\pi$-irrationality of $\alpha$ ensures that all eigenvalues $\lambda_{k}=e^{-ik\alpha}$ are distinct. Thus, the leading eigenvalues take the form
@@
Figure 6: Left: Eigenvalues of the Perron–Frobenius operator for a circle rotation by $\alpha=-1$ radians. Numerical eigenvalues (blue circles; leading pair corresponding to $k=\pm 1$ shown as red circles) coincide with the analytic spectrum (green crosses) Right: Magnitude versus argument of the same eigenvalues. The dashed green vertical lines mark angles that are negative integer multiples of $\alpha$, underscoring that the additional numerical eigenvalues are dynamically meaningful, but duplicate information from the lowest-order eigenvalues.
@@
We close this example by discussing the choice of $N=19$ basis functions. Recall we trained $\hat{\mathcal{L}}$ using trigonometric polynomials up to order 9 (equivalently linear combinations of real Fourier modes up to order 9). As discussed above, each two-dimensional real subspace spanned by $\theta\mapsto\sin(k\theta)$ and $\theta\mapsto\cos(k\theta)$ is an eigenspace with eigenpair $\lambda_{\pm k}$. Thus, in this special situation of circle rotation dynamics, the input and output data lie in the same 19-dimensional ($19=2\times 9+1$) subspace, which is itself invariant under $\mathcal{L}$. Consequently, SABON learns an invariant subspace in this example. Increasing the learned basis cardinality $N$ beyond 19 will not impact the prediction error, but the additional basis dimensions cannot easily be controlled with $9^{\rm th}$-order input/output data, and the resulting subspace will no longer be invariant.
3.4 Transfer operator for two nonlinear Anosov maps
We now examine two Anosov diffeomorphisms on the two-torus $\mathbb{T}^{2}$, both possessing well-defined SRB measures. The first is a weakly nonlinear perturbation of Arnold’s cat map, where the stable and unstable directions remain close to those of the original map. The second is obtained by conjugating the perturbed with a nonlinear diffeomorphism, yielding a strongly nonlinear map, which is far from volume preserving and whose stable and unstable directions vary significantly over the phase space
This pair of examples allows us to investigate how SABON performs as the geometric complexity of the dynamics increases. The strongly nonlinear conjugated map presents a more challenging test case, as the basis functions must adapt to curved unstable foliations and greater variation in local volume expansion and contraction.
3.4.1 A nonlinearly perturbed Arnold’s cat map